What volume (mL) of 0.025 M HNO3 is necessary to titrate 19.6 mL of 0.042 M Ca(OH)2 solution to the endpoint?



1
Expert's answer
2021-09-10T00:39:51-0400

The balanced reaction equation:

Ca(OH)2 + 2 HNO3 = Ca(NO3)2 + 2 H2O

The quantity of Ca(OH)2:

(0.042 M) * (0.01960 L) = 0.0008232 mol

So, the quantity of HNO3 is:

(0.0008232 mol) * 2 = 0.0016464 mol

Therefore, the volume must be:

(0.0016464 mol) / (0.025 M) = 0.065856 L = 65.856 mL

Answer: 65.856 mL